Please help.Question 10 0.9 pts As Sparkys Family Practice, M.D., maintains the accounting records of the clinic on a cash basis. DuringQuestion 11 1 pts Using the information presented above, determine Sparkys Total Operating Expenses on an accrual basis: $ (
Answer #1

Question 10

To calculate the service revenue on accrual basis, we need to use the below logic.

Service revenue (on accrual basis) = Cash collected from patients + Closing accounts receivable - Opening accounts receivable - Closing unearned revenue + Opening unearned revenue

Service revenue (on accrual basis) = 142600 + 15900 - 9200 + 2800 - 4100 = 148000

Question 11

Before considering the depreciation expenses and Bad debts expenses, we need to calculate the expenses accrued during the year 20X1.

Expenses on accrual basis = Expenses paid - Opening accrued expenses + closing accrued expenses + Opening prepaid expenses - Closing prepaid expenses

Expenses on accrual basis = 55470 - 3400 + 2300 + 1900 - 3400 = $52,870

Depreciation expenses on straight line basis = (Cost - Salvage value)/Useful life of the asset

Depreciation expenses = 40000/16 =2500

Allowance for doubtful debts on December 31, 20X1 = 2% of Accounts receivable = 2%×15,900 = 318

There is debit balance of $100 in allowance for doubtful debts account and we need the closing provision of $318. So, the bad debts expenses will be $418

Total operating expenses on accrual basis = Expense on accrual basis + depreciation expenses + bad debts expenses

Total operating expenses on accrual basis = 52870 + 2500 +418 = 55788